News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 24 May 2002 11.00am EDT

During the last week the observing schedule was interrupted following a high radiation event at 5:58am EDT on May 23 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and radiation levels dropped sufficiently to allow the loads to be restarted at 7.45am on May 24. A total of 93Ks of the observation of Sgr A* was impacted by the event.

A Chandra image release was made on May 22 showing a "true color" image of the supernova remnant N132D. The impressive image and further details about the SNR can be found at: http://chandra.harvard.edu/photo/2002/0050/index.html.

Of note this week were visits to the OCC by Dr. Jan Davis, Director, Flight Projects Directorate, MSFC and Mr. Bruce Dauer, Director of the Smithsonian Institution Office of Planning, Management and Budget, and next week by Mr. Alphonso Diaz, Director, GSFC.

The observing schedule for the next week is shown below and includes a continuation of the long observation of Sgr A* and an observation of AD Leo coordinated with the Hubble Space Telescope.

------------------------------------------
Radiation Belts                      May 27
Sgr A*                 ACIS-I        May 28
X1624-490              HRC-I         May 29
IM Nor                 ACIS-S
Radiation Belts                      May 30
V426 Oph               ACIS-S/HETG
NGC 4038/NGC 4039      ACIS-S        May 31
AD Leo                 ACIS-S/HETG   Jun 1
FTM0906+4952           ACIS-S
Radiation Belts
NGC 4459               ACIS-S        Jun 2
HS 1417+4722           ACIS-S
H162657                ACIS-I
M31 (5 obs)            HRC-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports